OpenStudy (anonymous):

If R is the total resistance for a parallel circuit with two resistors of resistances of r1 and r2 then 1/r=1/r1+1/r2 find the resistance r1 if the total resistance is 20 ohms and r2 is 75 ohms round your answer to the nearest ohms if neccesary a.16 ohms b.1405 ohms c.27 ohms d.102 ohms

OpenStudy (anonymous):

Just rearrange, plug, and chug:\[\frac{1}{r}=\frac{1}{r_1}+\frac{1}{r_2} \Longrightarrow \boxed{r_1=\dfrac{1}{\displaystyle \frac{1}{r}-\frac{1}{r_2}}}\]